# Mean-Force Hamiltonians from Influence Functionals

### v1

@@ Abstract
The Hamiltonian of mean force (HMF) provides the standard starting point for strong-coupling thermodynamics, yet explicit operator forms are known only in restricted settings. We present a “quenched density” framework that uses the Hubbard-Stratonovich transformation to rewrite the reduced equilibrium state as an average over local propagators in imaginary time. This approach rigorously separates the statistical definition of the environment from the algebraic structure of the system response. We apply this framework to the minimal case of a harmonic environment with a coupling commuting with the system Hamiltonian. In this scenario the correction to the HMF has an exact, closed-form expression. We validate this result against finite-bath trace-out calculations and stochastic imaginary-time sampling in a five-level projector-coupled model. Finally, we outline the program to extend this framework to non-Gaussian and non-commuting environments.
@@ I Introduction
What governs equilibrium? Historically, the canonical answer is the thermal Gibbs state, where the environment is characterised by a single parameter: the temperature. Yet, this canonical answer is puzzling. It implies that neither the specific character of the environment nor the form of its coupling to the system alters the system’s equilibrium. This of course is a consequence of the assumptions used to derive the answer - the weak-coupling or Markovian regimes open-systems theory has historically relied upon. At strong coupling however, the reduced state inherits explicit temperature dependence and interaction-induced operator content that definitionally cannot be captured by these assumptions [1, 2]. These issues have motivated a broad investigation into strong-coupling thermodynamics, including exact fluctuation relations [3, 4, 5], consistent heat and work definitions [6, 7, 8], and operational measurability [9, 10, 11].
Ultimately, while any subsystem in global equilibrium can formally be expressed in a canonical form, the effective Hamiltonian governing this state is generally not the bare Hamiltonian of the system. Instead, it must account for the non-trivial influence of the environment. If such an operator can be constructed, the standard description of equilibrium holds generically across all coupling strengths. The Hamiltonian of mean force (HMF) is the operator that reproduces this reduced equilibrium object as a Gibbs-like state, and provides the standard starting point for strong-coupling thermodynamics and mean-force Gibbs-state constructions [3, 12, 13, 14, 15].
@@
The ultimate target of thermodynamic descriptions is therefore unambiguous, and in fact a definition of HMF is equally unambiguous - it is simply the logarithm of the total state with the environment traced out. It is the representation of this operator that remains a challenge. Is the HMF a local operator? Does it admit a compact description in terms of few-body interactions? To address this, a veritable zoo of approximations - polaron transforms, weak-coupling expansions [16], reaction-coordinate mappings [17, 18] - have proliferated strong-coupling thermodynamics. These approximations often disagree in intermediate regimes however, and while powerful numerical techniques such as hierarchical equations of motion (HEOM) or path-integral Monte Carlo exist [19, 20, 21, 22, 23], they typically do not yield compact operator forms for the HMF.
@@
Ultimately, the source of these difficulties lie in the trace operation, which mixes the bath and system degrees of freedom in a way that is difficult to invert. For this reason, a general framework for representing the HMF has proved elusive. The purpose of this paper is to lay the groundwork for just such a framework. To do so, we construct a representation of the equilibrium density matrix that disentangles the statistical properties of the environment from the algebraic response of the system. We term this object the quenched density. By employing an imaginary-time influence functional [24, 25, 26] and its Hubbard-Stratonovich decoupling [27, 28, 29], we are able to rewrite the reduced equilibrium density not as a partial trace, but as a classical average over an ensemble of “quenched” imaginary-time propagators.
@@
In this picture, the bath is replaced by a stochastic field $\xi(\tau)$ [30, 31, 32], and the system evolves under a time-dependent effective Hamiltonian for each realization of the noise.
This formulation effectively sets the system in a frame where it can be expressed in terms of classical trajectories for the bath degrees of freedom. The Hubbard-Stratonovich transformation then allows us to exploit this to represent the bath information locally at the operator level. The framework thus provides a rigorous bridge between the nonlocal path-integral description (where the bath is exact but the system operator structure is opaque) and the local operator description required for thermodynamics.
We apply this framework to establishing a rigorous benchmark: the linearly coupled Gaussian bath in the commuting (pure-dephasing) sector [33, 34]. In this limit, relevant for quantum nondemolition measurements and qubit dephasing [35, 36], the technical obstruction of imaginary-time time-ordering vanishes, and the quenched density’s average can be performed analytically. We use this recover the known result that the bath induces a static potential renormalization $\lambda f^{2}$ (the reorganization energy) [37]. While this specific result is standard, its derivation within the quenched density framework serves two critical purposes. First, it demonstrates how the framework naturally recovers thermodynamic identities from path-integral primitives. Second, and most importantly, it provides a foundation for the generalisation of the framework to both non-commuting couplings and anharmonic environments.
The paper is structured as follows. In Sec. II we define the total Hamiltonian and reduced equilibrium objects. Sec. III develops the core of the framework: the Quenched Density representation of the imaginary-time propagator. Sec. IV connects this to the influence functional and derives the exact commuting Gaussian benchmark. Sec. V presents the numerical validation, confirming that the Quenched Density sampling, analytic prediction, and full finite-bath trace-out agree to machine precision. Sec. VI discusses how this Gaussian framework serves as the foundation for extending the theory to non-Gaussian and non-commuting environments.
@@ II The Total Hamiltonian and the Hamiltonian of Mean Force
Having specified the composite dynamics, we define the Hamiltonian of mean force by tracing the global Hamiltonian of mean force $H_{\mathrm{MF}}(\beta)$ as the operator whose Gibbs form reproduces the reduced equilibrium object up to a chosen normalisation. If the total system is at inverse temperature $\beta$, the unnormalised global equilibrium operator is $e^{-\beta H_{\mathrm{tot}}}$, then the reduced (unnormalise) equilibrium operator is
@@
For a harmonic bath linearly coupled through a collective coordinate, the bath can be eliminated exactly, yielding the Feynman-Vernon influence functional [24, 26, 2, 1]. In equilibrium this produces a Euclidean (imaginary-time) influence functional which is bilocal in $\tau$ and whose kernel is fixed by the bath spectral density $J(\omega)$ [19, 31, 39, 40]. While this gives an exact representation of the reduced equilibrium operator, it does not generically yield a compact local expression for $H_{\mathrm{MF}}$. The reduced equilibrium object is a path integral over the system degrees of freedom, rather than a simple exponential $e^{-\beta H_{\mathrm{eff}}}$ of a time-independent operator. Extracting $H_{\mathrm{MF}}$ - the logarithm of this ordered object - is therefore an inverse representability problem, not merely a question of coupling strength or temperature [12, 41].
@@ III The Quenched Density Representation
$Z_{\beta}\,\bar{\rho}(q_{f},q_{i};\beta)=\lim_{N\to\infty}\int\prod_{j=1}^{N}dq_{j}\prod_{j=0}^{N}\left\langle q_{j+1}\right|e^{-\Delta\bar{H}_{j}}\left|q_{j}\right\rangle.$ (14)
@@
$\bar{\rho}(q_{f},q_{i};\beta)=\lim_{\Delta\to 0}\frac{1}{Z_{\beta}}\int\prod_{j=1}^{N}dq_{j}\prod_{j=0}^{N}\left(\sqrt{\frac{m}{2\pi\Delta}}\right)\exp\!\left[-\sum_{j=0}^{N}\Delta\left(\frac{m}{2}\left(\frac{q_{j+1}-q_{j}}{\Delta}\right)^{2}+\bar{V}(q_{j},\tau_{j})\right)\right].$ (18)
In the limit $N\to\infty$ (i.e. $\Delta\to 0$), Eq. (18) becomes the Euclidean path integral [43, 44, 46]
@@ IV From Influence Functionals to the Hamiltonian of Mean Force
Having established the equivalence between the quenched density and the standard Euclidean path integral in the previous section, we now outline the specific form of the reduced system when coupled to a harmonic bath. Following the influence-functional formalism [24, 25, 26] and specifically its stochastic representation for Gaussian baths [31, 39, 40], the resulting reduced density matrix admits the imaginary-time path integral representation:
$\bar{\rho}_{Q}(q_{f},q_{i};\beta)=Z_{X}(\beta)\int_{q(0)=q_{i}}^{q(\beta)=q_{f}}\!\!\mathcal{D}q\;\exp\Bigg[-S_{Q}[q]+\frac{1}{2}\int_{0}^{\beta}d\tau\int_{0}^{\beta}d\tau^{\prime}f\big(q(\tau)\big)K(\tau-\tau^{\prime})f\big(q(\tau^{\prime})\big)\Bigg].$ (20)
@@
While the representation (20) is exact, it is nonlocal (bilocal) in imaginary time. This nonlocality can be resolved by a Hubbard - Stratonovich (HS) transformation [27, 28], which decouples the bilocal interaction by introducing an auxiliary Gaussian field $\xi(\tau)$. This step provides the bridge to an operator-level representation of $\bar{\rho}_{Q}$ as a Gaussian average over the local quenched canonical densities discussed in Sec. III:
$\begin{split}\bar{\rho}_{Q}(\beta)&=Z_{X}(\beta)\,\mathbb{E}_{\xi}\bigg[\\ &\quad\hat{\mathcal{T}}\exp\!\left(-\int_{0}^{\beta}d\tau\,\bar{H}(\tau;\xi)\right)\bigg],\end{split}$ (22)
The statistical properties of the auxiliary field $\xi(\tau)$ are set by the influence kernel [31, 39, 40]
@@
This operator representation maps directly to the quenched density framework. The challenge now lies in performing the average over $\xi$. To do so, we must isolate the stochastic term. Moving to the interaction picture with respect to $H_{Q}$, the propagator factorises as
@@
$\bar{\rho}_{Q}(\beta)=Z_{X}(\beta)\,e^{-\beta H_{Q}}\,\exp\!\left[\beta\lambda\,f^{2}\right].$ (28)
@@
$\mathrm{Tr}\,e^{-\beta H_{\mathrm{MF}}}=1$. The only nontrivial operator content is then the correction $\lambda f^{2}$, which is temperature independent.
A slight but interesting consequence of this result concerns integrable systems where the coupling operator $f$ is a conserved quantity ($H_{Q},f]=0$). In this case, the Hamiltonian of mean force $H_{\mathrm{MF}}$ shares the same eigenbasis as the bare system Hamiltonian. The strong coupling to the bath therefore does not mix the system energy levels, but rather induces a level-dependent energy shift $\lambda f_{n}^{2}$, where $f_{n}$ are the eigenvalues of the coupling operator. The integrability of the open system is thus preserved in the sense that the equilibrium state remains diagonal in the energy basis of the isolated system, but depending on the specific structure of $f$, the bath can induce a non-trivial operator-level effect on the reduced equilibrium state.
Interestingly $\lambda$ coincides with the static potential renormalisation $\frac{1}{2}\sum_{k}c_{k}^{2}/(m_{k}\omega_{k}^{2})$ of the Caldeira-Leggett model. If the standard counterterm $\lambda f^{2}$ has been absorbed into $H_{Q}$ (as stated in Sec. II), the bath-induced shift $\lambda f^{2}$ cancels it exactly. In this case, $H_{\mathrm{MF}}$ reduces to the bare system Hamiltonian $H_{Q}^{\mathrm{bare}}$, and the normalised mean-force Gibbs state coincides with the bare-system Gibbs state. In other words, in the commuting sector with translational invariance enforced, the bath has no operator-level effect on the reduced equilibrium state. Conversely, if no counterterm is included, $\lambda f^{2}$ represents a genuine bath-induced static renormalisation of the system potential. Amusingly, this result closely aligns to the master equation folk wisdom where superior results can be obtained when one neglects both counterterms and Lamb-shifts [49].

### v2

@@ Abstract
The Hamiltonian of mean force (HMF) provides the standard starting point for strong-coupling thermodynamics, yet explicit operator forms are known only in restricted settings. We present a “quenched density” framework that uses the Hubbard-Stratonovich transformation to rewrite the reduced equilibrium state as an average over local propagators in imaginary time. This approach rigorously separates the statistical definition of the environment from the algebraic structure of the system response. We apply this framework to the minimal case of a harmonic environment with a coupling commuting with the system Hamiltonian. In this scenario the correction to the HMF has an exact, closed-form expression. We validate this result against finite-bath trace-out calculations and stochastic imaginary-time sampling in a five-level projector-coupled model. This confirms the specific strong-coupling projector dominance predicted by energetic arguments
@@ I Introduction
What governs equilibrium? The canonical answer is the thermal Gibbs state, where the environment is characterised by a single parameter: the temperature. Yet, this canonical answer is puzzling. It implies that neither the specific character of the environment nor the form of its coupling to the system alters the system’s equilibrium. This of course is a consequence of the assumptions used to derive the answer - the weak-coupling or Markovian regimes open-systems theory has historically relied upon [1]. At strong coupling however, the reduced state inherits explicit temperature dependence and interaction-induced operator content that definitionally cannot be captured by these assumptions [2, 3]. These issues have motivated a broad investigation into strong-coupling thermodynamics, including exact fluctuation relations [4, 5, 6], consistent heat and work definitions [7, 8, 9], and operational measurability [10, 11, 12]. This representability question sits alongside operational ambiguities in thermodynamic bookkeeping beyond weak coupling, where heat/work/state assignments can depend on the chosen framework [13].
While formally one can write the reduced state in canonical exponential form, the effective Hamiltonian governing this state is generally not the bare Hamiltonian of the system. Instead, it must account for the non-trivial influence of the environment in a manner that goes beyond the specification of temperature. If an effective operator capturing this influence can be constructed, the standard description of equilibrium holds generically across all coupling strengths. The Hamiltonian of mean force (HMF) is the operator that reproduces this reduced equilibrium object as a Gibbs-like state, and provides the standard starting point for strong-coupling thermodynamics[4, 5, 14, 15, 16]. Historically, much of open-system theory emphasised weak-coupling and Markovian regimes, where reduced equilibrium states are well approximated by Gibbs states of renormalized system Hamiltonians. At finite coupling, equilibrium consistency is anchored in the correlated global Gibbs state, and the reduced operator inherits coupling-dependent structure that need not be captured by a simple renormalization. The HMF can become temperature dependent and can encode interaction-induced terms that are absent in $H_{Q}$, including effective many-body or nonlocal operator content[15, 16, 7, 2, 3, 17]. This is the conceptual tension: the reduced equilibrium state is well defined, but its generator is not generally simple. In the strong-coupling regime one instead characterises equilibrium through the mean-force Gibbs state, whose structure and limiting forms (including ultrastrong-coupling projector dominance) have been analysed from a ‘state-first’ perspective [18, 19]. The present work is complementary: we start from the Euclidean influence functional and give an explicit local stochastic representation that yields a closed operator generator in a certified commuting–Gaussian sector. Related approaches connecting the HMF and influence functional have recently been explored in Refs. [20, 21].
@@
The ultimate target of thermodynamic descriptions is therefore unambiguous, and in fact a formal definition of HMF is equally unambiguous - it is simply the logarithm of the total state after tracing out its environment. It is the representation of this operator that remains a challenge. Is the HMF a local operator? Does it admit a compact description in terms of few-body interactions? To address this, a veritable zoo of approximations - polaron transforms [22], weak-coupling expansions [18], reaction-coordinate mappings  [23, 24, 25, 26, 27] - have proliferated strong-coupling thermodynamics. These approximations often disagree in intermediate regimes however, and while powerful numerical techniques have been derived from path-integrals [28, 29, 30, 31, 32, 33], they typically do not yield compact operator forms for the HMF.
@@
Ultimately, the source of these difficulties lie in the trace operation, which mixes the bath and system degrees of freedom in a way that is difficult to invert. For this reason, a general framework for representing the HMF has proved elusive. The purpose of this paper is to lay the groundwork for just such a framework. To do so, we construct a representation of the equilibrium density matrix that disentangles the statistical properties of the environment from the algebraic response of the system. We term this object the quenched density, and to obtain it we employ the influence functional approach [34, 35, 36]. This formalism provides a natural language for the problem. For a recent unifying overview that links hierarchical, chain-mapping, pseudomode, stochastic, and extended-state-space approaches to non-Markovian open-system dynamics, see Ref. [37]. By employing the influence functional in concert with the Hubbard-Stratonovich decoupling [38, 39, 40], we are able to rewrite the reduced equilibrium density not as a partial trace, but as a classical average over an ensemble of “quenched” imaginary-time propagators.
@@
In this picture, the bath is replaced by a stochastic field $\xi(\tau)$ [41, 42, 43, 44, 45, 46], and the system evolves under a time-dependent effective Hamiltonian for each realisation of the noise.
This formulation is obtained in two steps. The path integral sets the system in a frame where it can be expressed in terms of classical trajectories for the bath degrees of freedom, which can be integrated out. The resulting action is then a non-local kernel in imaginary time. The Hubbard-Stratonovich transformation remedies this by representing the bath information locally. This then allows the path integral to be reversed and the reduced density to be formulated at the operator level. The framework thus provides a rigorous bridge between the nonlocal path-integral description (where the bath is exact but the system operator structure is opaque) and the local operator description required for thermodynamics.
We apply this framework to establishing a rigorous benchmark: the linearly coupled Gaussian bath in the commuting (pure-dephasing) sector [47, 1]. In this limit, relevant for quantum nondemolition measurements and qubit dephasing [48, 49], the technical obstruction of imaginary time ordering vanishes, and the quenched density’s average can be performed analytically. We use this recover the known result that the bath induces a static potential renormalisation $\lambda f^{2}$ (the reorganisation energy) [50]. While this specific result is standard, its derivation within the quenched density framework serves two critical purposes. First, it demonstrates how the framework naturally recovers thermodynamic identities from path-integral primitives. Second, and most importantly, it provides a foundation for the generalisation of the framework to both non-commuting couplings and anharmonic environments.
The paper is structured as follows. In Sec. II we define the total Hamiltonian and reduced equilibrium objects. Sec. III develops the core of the framework: the Quenched Density representation of the imaginary-time propagator. Sec. IV connects this to the exact commuting Gaussian benchmark. Sec. V presents the numerical validation, confirming that the quenched density sampling, analytic prediction, and full finite-bath trace-out agree to machine precision. Sec. VI closes the paper with a summary of results, and highlights the essential contribution of this framework as a foundation for the future extension of the theory to non-Gaussian and non-commuting environments.
@@ II The Total Hamiltonian and the Hamiltonian of Mean Force
Having specified the composite dynamics, we define the Hamiltonian of mean force by tracing the global Hamiltonian of mean force $H_{\mathrm{MF}}(\beta)$ as the operator whose Gibbs form reproduces the reduced equilibrium object up to a chosen normalisation. If the total system is at inverse temperature $\beta$, the unnormalised global equilibrium operator is $e^{-\beta H_{\mathrm{tot}}}$, then the reduced (unnormalised) equilibrium operator is
@@
For a harmonic bath linearly coupled through a collective coordinate, the bath can be eliminated exactly, yielding the Feynman-Vernon influence functional [34, 36, 3, 2]. In equilibrium this produces a Euclidean (imaginary-time) influence functional which is bilocal in $\tau$ and whose kernel is fixed by the bath spectral density $J(\omega)$ [28, 42, 52, 53]. While this gives an exact representation of the reduced equilibrium operator, it does not generically yield a compact local expression for $H_{\mathrm{MF}}$. The reduced equilibrium object is a path integral over the system degrees of freedom, rather than a simple exponential $e^{-\beta H_{\mathrm{eff}}}$ of a time-independent operator. Extracting $H_{\mathrm{MF}}$ - the logarithm of this ordered object - is therefore an inverse representability problem [54, 55, 56, 57], not merely a question of coupling strength or temperature [15, 58].
@@ III The Quenched Density Representation
In limiting regimes the same Euclidean influence-functional machinery collapses to explicitly local reduced descriptions, e.g. in the strong-friction quantum Smoluchowski limit. This provides a useful bridge: locality can emerge either by algebraic closure (our route) or by a friction-dominated dynamical limit [62].
@@
$\bar{\rho}(q_{f},q_{i};\beta)=\lim_{N\to\infty}\int\prod_{j=1}^{N}dq_{j}\prod_{j=0}^{N}\left\langle q_{j+1}\right|e^{-\Delta\bar{H}_{j}}\left|q_{j}\right\rangle.$ (14)
@@
$\bar{\rho}(q_{f},q_{i};\beta)=\lim_{\Delta\to 0}\int\prod_{j=1}^{N}dq_{j}\prod_{j=0}^{N}\left(\sqrt{\frac{m}{2\pi\Delta}}\right)\exp\!\left[-\sum_{j=0}^{N}\Delta\left(\frac{m}{2}\left(\frac{q_{j+1}-q_{j}}{\Delta}\right)^{2}+\bar{V}(q_{j},\tau_{j})\right)\right].$ (18)
In the limit $N\to\infty$ (i.e. $\Delta\to 0$), Eq. (18) becomes the Euclidean path integral [60, 61, 64], such that the reduced density (normnalised by $Z_{\bar{H}}$) is given by
@@ IV From Influence Functionals to the Hamiltonian of Mean Force
Having established the equivalence between the quenched density and the standard Euclidean path integral in the previous section, we now outline the specific form of the reduced system when coupled to a harmonic bath. Following the influence-functional formalism [34, 35, 36] and specifically its stochastic representation for Gaussian baths [42, 52, 53]. The resulting reduced density matrix admits the imaginary-time path integral representation:
$\bar{\rho}_{Q}(q_{f},q_{i};\beta)=\int_{q(0)=q_{i}}^{q(\beta)=q_{f}}\!\!\mathcal{D}q\;\exp\Bigg[-S_{Q}[q]+\frac{1}{2}\int_{0}^{\beta}d\tau\int_{0}^{\beta}d\tau^{\prime}f\big(q(\tau)\big)K(\tau-\tau^{\prime})f\big(q(\tau^{\prime})\big)\Bigg].$ (20)
@@
Related exact Gaussian-bath embeddings that trade nonlocal memory for time-local dynamics in a minimally extended state space are developed in Ref. [67]; here we instead emphasise the Euclidean equilibrium generator and its exact closure conditions.
While the representation in Eq.(20) is exact, it is nonlocal (bilocal) in imaginary time. This nonlocality can be resolved by a Hubbard - Stratonovich (HS) transformation [38, 39], which decouples the bilocal interaction by introducing an auxiliary Gaussian field $\xi(\tau)$. This step provides the bridge to an operator-level representation of $\bar{\rho}_{Q}$ as a Gaussian average over the local quenched canonical densities discussed in Sec. III:
$\bar{\rho}_{Q}(\beta)=\mathbb{E}_{\xi}\bigg[\hat{\mathcal{T}}\exp\!\left(-\int_{0}^{\beta}d\tau\,\bar{H}(\tau;\xi)\right)\bigg],$ (22)
where $\bar{H}(\tau;\xi)=H_{Q}-\xi(\tau)f$ is the time-dependent quenched Hamiltonian. The statistical properties of the auxiliary field $\xi(\tau)$ are set by the influence kernel [42, 52, 53]
@@
This Hubbard–Stratonovich unravelling is the same as that underlying partition-free stochastic Liouville formulations of Gaussian environments and equilibrated initial conditions [42]. Recent work also develops a time-local minimally extended-state-space embedding (QD-MESS) via influence-functional decomposition and Hubbard–Stratonovich unravelling, recovering Caldeira–Leggett and Smoluchowski-type limits as special cases [67].
Having performed this transformation, the operator representation maps directly to the quenched density framework. The challenge now lies in performing the average over $\xi$. To do so, we must isolate the stochastic term. Moving to the interaction picture with respect to $H_{Q}$, the propagator factorises as
@@
$\bar{\rho}_{Q}(\beta)=e^{-\beta H_{Q}}\,\exp\!\left[\beta\lambda\,f^{2}\right].$ (28)
@@
The only nontrivial operator content is the correction $\lambda f^{2}$, which is temperature independent.
An immediate consequence that can be inferred from this result is that it is identical to the result in the classical mean-force construction [50, 15, 68]. This equilvance can be understand mostly easily through the link to Koopman von-Neumann dynamics. In real time, classical evolution is generated by the Liouville/Koopman operator $\mathcal{L}_{H}=\{\,\cdot\,,H\}$, with propagator $e^{t\mathcal{L}_{H}}$. To obtain Gibbs states from imaginary time however, the relevant generator becomes the Hamiltonian. This is true in both the quantum and classical scenarios, but in the latter case all observable operators (and hence the Hamiltonian) commute identically. The classical equilibrium construction will therefore always lie within the commuting sector of the fully quantum problem. For a fuller discussion of the dynamical distinctions between quantum and classical systems in both real and imaginary time, see ref. [69].
Another slight (but interesting) consequence of this result concerns integrable systems where the coupling operator $f$ is a conserved quantity ($H_{Q},f]=0$). In this case, the Hamiltonian of mean force $H_{\mathrm{MF}}$ shares the same eigenbasis as the bare system Hamiltonian. The strong coupling to the bath therefore does not mix the system energy levels, but rather induces a level-dependent energy shift $\lambda f_{n}^{2}$, where $f_{n}$ are the eigenvalues of the coupling operator. The integrability of the open system is thus preserved in the sense that the equilibrium state remains diagonal in the energy basis of the isolated system, but depending on the specific structure of $f$, the bath can induce a non-trivial operator-level effect on the reduced equilibrium state.
A further important consequence is that $\lambda$ coincides with the static potential renormalisation $\frac{1}{2}\sum_{k}c_{k}^{2}/(m_{k}\omega_{k}^{2})$ of the Caldeira-Leggett model. If the standard counterterm $\lambda f^{2}$ has been absorbed into $H_{Q}$ (as stated in Sec. II), the bath-induced shift $\lambda f^{2}$ cancels it exactly. In this case, $H_{\mathrm{MF}}$ reduces to the bare system Hamiltonian $H_{Q}^{\mathrm{bare}}$, and the normalised mean-force Gibbs state coincides with the bare-system Gibbs state. In other words, in the commuting sector with translational invariance enforced, the bath has no operator-level effect on the reduced equilibrium state. Conversely, if no counterterm is included, $\lambda f^{2}$ represents a genuine bath-induced static renormalisation of the system potential. Amusingly, this result closely aligns to the master equation folk wisdom where superior results can be obtained when one neglects both counterterms and Lamb-shifts [17].